40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4515     Entry time: Tue Apr 12 12:01:30 2011
Author: josephb 
Type: Update 
Category: General 
Subject: IFO controls, now with 10% less lying (ITMX/ITMY controls swapped) 

The ITMX/ITMY control swap is complete.

The steps from this elog were followed.

In addition, I did a burt restore of c1sus, c1mcs.

I then swapped all the gain settings from ITMX to ITMY, and reenabled the watchdogs.

I did some basic kick tests (1000 counts into UL coil) and confirmed channels like C1:SUS-ITMX_ULPD_VAR (watchdogs mV readback) corresponded to the correct optic.  I also checked that the POS, PIT, YAW, SIDE produced reasonable damping when engaged.

ELOG V3.1.3-